### v1

@@ 3 ML Model
    VM lifetimes depend on many factors that cannot be described by its features (e.g., human behavior or external events). As such, some VMs are fundamentally impossible to predict. Predicting a single VM lifetime thus limits accuracy as shown in Figure 2.
@@
Our model addresses these limitations by predicting a probability distribution as opposed to a single value. This approach is well-known in survival analysis Wang et al. (2019). Survival models capture the “uncertainty” associated with a particular set of feature values. 1) If a VM is highly predictable based on the features, the distribution will be narrow, while difficult-to-predict VMs have wider distributions. 2) We use distributions to update our VM estimates based on the lifetime so far (uptime). Instead of relying on the initial prediction, we compute the conditional expectation $\mathbb{E}(T_{r}|T_{u})$ for an updated estimate of a VM’s lifetime over time: “Given a VM has been running for interval $T_{u}$, what is the expected remaining lifetime $T_{r}$?”. This value can be directly calculated from the PDF in Figure 2.
This approach avails itself to a range of models, including different forms of survival models, neural networks, and gradient-boosted decision trees (GBDTs). GBDTs performed best in our experiments. Since they are not distribution-based, we pass an additional parameter $T_{u}$ into the model that represents the uptime of the VM so far (i.e., the x axis of the PDF, where lifetime = uptime + remaining lifetime), and predict the remaining lifetime. We augment every training example by turning it into multiple examples with different values of $T_{u}$: specifically, 12.5%, 25%, … of the original lifetime. This implementation turns a regression model into a survival model with good accuracy. We train the model directly on $\mathbb{E}(T_{r}|T_{u})$ as the label.
@@
Our models use a range of features, including the zone, VM family, VM shape, whether or not the VM attaches to local SSD, and internal metadata. We use a production-grade machine learning library called Yggdrasil Guillame-Bert et al. (2023). We generate our training data by querying a large-scale internal data warehouse that contains historical data of VMs. We use a combination of distributed SQL queries and data processing pipelines to generate a set of labeled examples and feed them into the training framework. Our data sets are on the order of 1M examples for training. The resulting model is loaded from a file system and runs in binaries via a range of language bindings. We periodically retrain our model on recent data.
@@
An important design consideration is how to deploy the models. For instance, Barbalho et al. Barbalho et al. (2023) run their models on separate inference servers. They cache prediction results to tolerate unavailability of servers, increasing complexity. Running model inference on separate servers also requires handling model rollouts and verification separately since they are not integrated with existing rollout mechanisms.
We choose instead to embed the model directly into the Borg binary. This deployment reduces possible failure modes. By piggybacking on Borg, we update models with the same frequency as Borg, testing the two together. This approach lowers model overhead and achieves a median latency of 9 us (Figure 8), which is 780$\times$ lower than the median latency for LA Barbalho et al. (2023). With this low latency, we can run our model more frequently, including for lifetime-guided defragmentation (Section 4.4) and to update and correct mispredictions.
Our model has been deployed in production for over 9 months and has maintained all our production standards, while improving empty machines, reducing stranding, and reducing VM migrations.
@@ 4 Scheduling Algorithms
We show that repredicting lifetimes offers a fundamental advantage over one-shot prediction in our multi-dimensional scheduling approach. We introduce three algorithms. LAS is non-invasive and becomes active when a set of hosts satisfy all other criteria and thus have the same score without the bin packing score (Section 4.2). LAVA, in contrast, makes lifetime-aware scheduling a fundamental part of the scheduler (Section 4.3). LARS exploits lifetimes for defragmentation and maintenance (Section 4.4).
@@ 4.1 Theoretical Foundation
We can show analytically that repredicting lifetimes and correcting for mispredictions fundamentally improves how well any given algorithm can do. Specifically, if the initial error in lifetime prediction is a positive constant, then the number of hosts required without correcting for mispredictions will exceed the number for the same best fit algorithm with repredictions by $\Omega(m)$, where $m$ is the number of hosts.
4.2 LAS: Non-Invasive Lifetime-Aware Scheduling
The goal of LAS (Figure 4) is to create more empty machines by scheduling VMs on a host where all other VMs are likely to exit at a similar time or later. Given a VM, LAS ranks potential hosts based on the exit times for the VMs already scheduled on that host. It computes the maximum of the predicted exit time of all VMs on the host. This score is conceptually similar to LA Barbalho et al. (2023), with a crucial difference. Instead of using the original VM lifetime predictions, LAS repredicts the remaining lifetime (i.e., the updated exit time) for all VMs, using their uptime so far.
LAS corrects any earlier mispredictions as follows. Consider a host where all remaining VMs have exceeded their originally predicted lifetime. Without correcting mispredictions, the scheduler has to assume that this host will soon become empty and will avoid scheduling long-lived VMs on this host. We hypothesize that this choice is the reason why LA is sensitive to prediction accuracy, leading the authors to not deploy LA. In contrast, LAS schedules new VMs on hosts using the repredicted remaining lifetimes of all VMs on each host. As such, LAS continues to improve its predictions if the original prediction was wrong.
Figure 4: Overview of the LAS algorithm.
@@ 4.2 LAS: Non-Invasive Lifetime-Aware Scheduling
We add the temporal cost into our lexicographic scoring function one level above the bin packing score. The temporal cost never has an effect on any higher-ranked business metrics. It only becomes the deciding factor if all those scores are the same. The score is thus non-invasive. It is entirely local to the bin packing score and is equivalent to modifying the bin packing score to take a third dimension (lifetime) into account.
4.3 LAVA
A limitation of LAS is that mispredictions may gradually bump up the host lifetime. By trying to match the new VM to the exit times of other VMs on a host, the host may never free up because each under-predicted VM further bumps up the exit time, causing even more VMs to be scheduled, which in turn bumps up the exit time further.
@@ 4.3 LAVA
LAVA divides lifetime predictions into lifetime classes: $<$1h, 1-10h, 10-100h, 100-1000h. We refer to these lifetime classes as LC1, LC2, L3, and L4. Each host is also assigned a lifetime class. LAVA runs an additional coarse-grained host scoring function based on their lifetime classes and breaks ties with LAS.
Each host has a lifetime class and is in one of three states: empty, open, or recycling (similar to LLAMA Maas et al. (2020)). When a new VM arrives and there is no eligible open or recycling host, LAVA places the VM on an empty host, assigns the host the open state, and assigns the host the lifetime class of the VM. As long as this host is open, LAVA will only schedule other VMs of the same lifetime class on it (Figure 5(a)). Once over 90% of the resources (CPU or memory) of an open host are occupied, LAVA transitions the host to the recycling state. All the VMs that were present on the host when it transitioned state are labeled residual VMs (i.e., the VMs of the original lifetime class of the host). Recycling hosts only receive VMs that are at least one lifetime class lower, e.g., LAVA only places new VMs with a predicted lifetime of 10 hours or shorter on 100-hour recycling hosts.
Building upon LAS, LAVA is integrated into the existing hierarchical scheduler by adding another scoring function before LAS. This higher-ranked scoring function keeps track of additional states such as predicted host lifetime class, host state, and deadline for each host. The size of this additional state is negligible compared to the substantial amount of existing per-VM state.
When a new VM request arrives, LAVA scores the hosts in the following order (in decreasing preference): recycling hosts whose lifetime classes are greater than the VM lifetime class (the closer the more preferred), open hosts whose lifetime class are equal to the VM lifetime class, any non-empty host, and lastly empty hosts. In contrast to LAS, which acts as a tie-breaker for the existing algorithm, LAVA makes decisions that the baseline could not have made, and is therefore invasive.
@@ 4.4 LARS: Defragmentation & Maintenance
The above algorithms focus on reducing fragmentation by improving VM placement. Borg also performs software and hardware maintenance activities, applying security, microcode, host OS, and VM software updates. It also actively defragments hosts for some VM families. These activities live migrate VMs. Live migration uses pre- and post-copy techniques that minimize VM execution time disruptions to seconds. However, total migration times are proportional to VM memory and SSD usage, and consume capacity on both hosts. In simulations, we model that both hosts are busy for a conservative 20 minutes.
Borg reserves a small fraction of hosts for defragmentation, maintenance, and to tolerate hardware failures, and thus these activities do not impact available capacity. We show how using lifetimes can reduce the number of VMs that are disrupted during defragmentation and maintenance. We call the algorithm LARS. The rest of our exposition focuses on defragmentation for conciseness.
@@
Note that LARS uses the same algorithm to choose a target host for migration as it does during initial VM placement. As such, LAS and LAVA use repredictions from our distribution-based model to place the VM on a better-suited host than it was placed initially during migration.
@@ 5 Implementation
Figure 6: Empty host improvements comparing LAS, LAVA with LA-Binary, to our production baseline for 24 C2 pools in simulation. On average, LAVA increases empty hosts by 6.5 pp, compared to 6.1 pp for LAS, and 5.0 pp for LA-Binary.
Our algorithms are integrated into Borg’s main scheduler code base. The scheduler runs within Borg Prime, which handles incoming VM requests and assigns them to hosts. The Borg Prime binary is replicated across a number of instances, with one of them serving as the elected leader. All replicas run the same binary and consensus is handled at a lower layer. From the perspective of our algorithms, we treat this scenarios as if there was a single Borg Prime instance.
As mentioned in Section 3, our model is compiled into the Borg Prime binary and rolled out with Borg. This coupling has the advantage that canary deployments that are part of our rollout process test the entire end-to-end approach. We deploy our software regularly, and may support updating our model as often as once a week. However, we found that the models’ accuracy remains high even a month later as shown in Figure 10. Note that while we found it highly advantageous from a latency and complexity perspective to compile the model into the binary, our approach is not restricted to this deployment scenario.
We carefully measured the CPU and DRAM overheads of our approach in production. Binary size increases by less than 10 MB, and both the memory and CPU consumption of our approach are indistinguishable from noise. This lack of observable overhead may seem counter-intuitive, but a VM record has so much state and Borg performs so much computation during scoring that the extra bytes of memory and the 9 us of model latency are not visible. We process on the order of 10-100 scheduling requests per second in each cluster, and even hypothetically repredicting 100 VMs in the process would add less than 1 ms. In practice, we repredict much fewer VMs, since most hosts are filtered out by higher-ranking scoring functions and only VMs on considered hosts are re-scored. The computation is parallelized as well.
@@ 5.1 Simulations
LAS/LAVA: We developed an event-driven simulator to explore our approach. We extract production traces of VM start, exit, and restart events within a particular pool and then replay this trace against a simulated instance of the scheduler. The simulated scheduler uses our production code base and is thus highly accurate. We validated our simulator against production.
LARS: To calculate the benefit of LARS (lifetime-based defragmentation), we built a modified simulator, using the same production traces. From our traces, we collect the full list of migrations that were performed during a time interval. In production, we limit the number of live migrations that can be in progress at a time to batches of 3. We simulate this behavior by assuming that all migrations are performed in a certain order (in our baseline, defined by the trace), but have to wait until a slot is available. This approach has the effect that some VMs exit while others are migrating. LARS modifies this order based on lifetime predictions.
@@ 5.2 Production Measurements
Measuring the impact of scheduling policies in production is challenging. Since workloads and total capacity in a pool are constantly changing, it is difficult to attribute metric changes to a particular modification. We therefore create an A/B testing methodology. We split the hosts into two halves and apply our scheduling algorithm to one of them. Our production metrics report empty hosts, stranding, and model prediction for the entire pool. However, stranding is only accurate for entire pools and cannot differentiate A/B hosts. We report live production experiments in four pools, where we apply the algorithm to the entire pool and evaluate scheduling latency. These experiments also ensure that our A/B approach does not distort behavior. Note that these A/B deployments are at production scale. The cluster/zone sizes for this product range from 100s to 10,000+ hosts with O(100-10,000) VMs active at a time.
LAS has been deployed fleet-wide for one VM family for months, at the time of writing. Time series data from the fleet-wide production rollout confirmed our simulation numbers and A/B experiments, and show a fleet-wide increase of empty hosts by 4 pp for this VM family.
@@ 5.3 LA-Binary Comparison
Our main comparison is to LA-Binary Barbalho et al. (2023). Since our data center environment is different, perfectly replicating their setup is impossible, e.g., the ML model features are different. We are faithful to LA-Binary’s description and favorable to LA-Binary when in doubt. For an apples-to-apples comparison, we use the exact same model when comparing LA-Binary to LAS and LAVA, but without repredictions for LA-Binary. VMs are classified as short or long lived based on their lifetime predictions with the same two-hour cutoff threshold as in their paper.

### v2

@@ 3 ML Lifetime Prediction Model
    VM lifetimes depend on many factors that cannot be described by its features, e.g., human behavior or external events. As such, some VMs are fundamentally impossible to predict. Predicting a single VM lifetime thus limits accuracy, as shown in Figure 2.
@@
Our model addresses these limitations by predicting a probability distribution as opposed to a single value. This approach is well-known in survival analysis Wang et al. (2019). Survival models capture the “uncertainty” associated with a particular set of feature values. If a VM is highly predictable based on the features, the distribution will be narrow, while difficult-to-predict VMs have wider distributions. Here, we use distributions to update our VM estimates based on the lifetime so far (uptime). Instead of relying on the initial prediction, we compute the conditional expectation $\mathbb{E}(T_{r}|T_{u})$ for an updated estimate of a VM’s lifetime over time: “Given a VM has been running for interval $T_{u}$, what is the expected remaining lifetime $T_{r}$?”. This value is directly calculated from the PDF in Figure 2.
This approach avails itself to a range of models and we compare a few (Appendix B), including different forms of survival models, neural networks, and gradient-boosted decision trees (GBDTs). GBDTs performed best in our experiments. Since they are not distribution-based, we pass an additional parameter $T_{u}$ into the model that represents the uptime of the VM so far (i.e., the x axis of the PDF), and we predict the remaining lifetime $T_{r}$. We augment every training example by turning it into multiple examples with different values of $T_{u}$: specifically, 12.5%, 25%, … of the original lifetime. This implementation turns a regression model into a survival model with good accuracy. We train the model directly on $\mathbb{E}(T_{r}|T_{u})$ as the label.
@@
Our models use a range of features, including the zone, VM family, VM shape, whether or not the VM attaches to local SSD, and internal project metadata (Appendix A). We use a production-grade machine learning library called Yggdrasil Guillame-Bert et al. (2023). We generate our training data by querying a large-scale internal data warehouse that contains historical data of VMs. We use a combination of distributed SQL queries and data processing pipelines to generate a set of labeled examples and feed them into the training framework. Our data sets are on the order of 1M examples for training. The resulting model is loaded from a file system and runs in binaries via a range of language bindings. We periodically retrain our model on recent data.
@@
An important design consideration is how to deploy the models. For instance, Barbalho et al. Barbalho et al. (2023) run their models on separate inference servers. They cache prediction results to tolerate unavailability of servers, increasing complexity. Running model inference on separate servers also requires handling model rollouts and verification separately since they are not integrated with existing rollout mechanisms. We choose instead to embed the model directly into the Borg binary. This deployment reduces possible failure modes. By piggybacking on Borg rollouts, we update models with the same frequency as Borg, testing the two together. This approach lowers model overhead and achieves a median latency of 9 us (Figure 9), which is 780$\times$ lower than the median latency for LA Barbalho et al. (2023). With this low latency, we can run our model more frequently, including for lifetime-guided defragmentation (Section 4.4) and to update and correct mispredictions.
We deployed our model in production over a year ago. It maintained all our production standards, while improving empty machines, stranding, and VM migrations.
@@ 4 Scheduling Algorithms
We show that repredicting lifetimes offers a fundamental advantage over one-shot prediction in our multi-dimensional scheduling approach. We introduce three algorithms. NILAS is non-invasive and becomes active when a set of hosts satisfy all other criteria and thus have the same score without the bin packing score (Section 4.2). LAVA, in contrast, makes lifetime-aware scheduling a fundamental part of the scheduler (Section 4.3). Note that while LA and NILAS place VMs with similar lifetime together, the key idea of LAVA is the opposite – it adds many short lived VMs to hosts with one or more long lived VM. The intent is to create hosts that have a wider range of different lifetimes, with an upper bound that is decreasing over time. LARS exploits lifetimes to reduce VM migrations during defragmentation and maintenance (Section 4.4).
@@ 4.1 Theoretical Foundation
We show analytically that repredicting lifetimes and correcting for mispredictions fundamentally improves how well any algorithm can do. Specifically, if the initial error in lifetime prediction is a positive constant, then the number of hosts required without correcting for mispredictions will exceed the number for the same best fit algorithm with repredictions by $\Omega(m)$, where $m$ is the number of hosts. Appendix E includes the precise theorem and proof.
4.2 NILAS: Non-Invasive Lifetime-Aware Scheduling
The goal of NILAS (Figure 4) is to create more empty machines by scheduling VMs on a host where all other VMs are likely to exit at a similar time or later. Given a VM, NILAS ranks potential hosts based on the exit times for the VMs already scheduled on that host. It computes the maximum of the predicted exit time of all VMs on the host. This score is conceptually similar to LA Barbalho et al. (2023), with a crucial difference: Instead of using the original VM lifetime predictions, NILAS repredicts the remaining lifetime (i.e., the updated exit time) for all VMs, using their uptime so far.
NILAS corrects any earlier mispredictions as follows. Consider a host where all remaining VMs have exceeded their originally predicted lifetime. Without correcting mispredictions, the scheduler has to assume that this host will soon become empty and will avoid scheduling long-lived VMs on this host. We hypothesize that this choice is the reason why LA is sensitive to prediction accuracy, leading the authors to not deploy LA. In contrast, NILAS schedules new VMs on hosts using the repredicted remaining lifetimes of all VMs on each host. As such, NILAS continues to improve its predictions if the original prediction was wrong.
Figure 4: Overview of the NILAS algorithm.
@@ 4.2 NILAS: Non-Invasive Lifetime-Aware Scheduling
We add the temporal cost into our lexicographic scoring function one level above the bin packing score. The temporal cost never has an effect on any higher-ranked business metrics. It only becomes the deciding factor if all those scores are the same. The score is thus non-invasive. It is local to the bin packing score and is equivalent to modifying the bin packing score to add a third dimension (lifetime).
4.3 LAVA: Lifetime-Aware VM Allocation
A limitation of NILAS is that mispredictions may gradually bump up the host lifetime. By trying to match the new VM to the exit times of other VMs on a host, the host may never free up because each under-predicted VM further bumps up the exit time, causing even more VMs to be scheduled, which in turn bumps up the exit time further.
@@ 4.3 LAVA: Lifetime-Aware VM Allocation
LAVA divides lifetime predictions into lifetime classes: $<$1h, 1-10h, 10-100h, 100-1000h. We refer to these lifetime classes as LC1, LC2, L3, and L4. Each host is also assigned a lifetime class. LAVA runs an additional coarse-grained host scoring function based on their lifetime classes and breaks ties with NILAS.
@@
Each host has a lifetime class and is in one of three states: empty, open, or recycling, similar to LLAMA Maas et al. (2020). When a new VM arrives and there is no eligible open or recycling host, LAVA places the VM on an empty host, assigns the host the open state, and assigns the host the lifetime class of the VM. As long as this host is open, LAVA will only schedule other VMs of the same lifetime class on it (Figure 5(a)). Once over 90% of the resources (CPU or memory) of an open host are occupied, LAVA transitions the host to the recycling state. All the VMs that were present on the host when it transitioned state are labeled residual VMs (i.e., the VMs of the original lifetime class of the host). Recycling hosts only receive VMs that are at least one lifetime class lower, e.g., LAVA only places new VMs with a predicted lifetime of 10 hours or shorter on 100-hour recycling hosts.
LAVA is also integrated into Borg’s existing hierarchical scheduler, but adds another scoring function before NILAS. This higher-ranked scoring function keeps track of additional states such as predicted host lifetime class, host state, and deadline for each host. The size of this additional state is negligible compared to other per-VM state.
When a new VM request arrives, LAVA scores the hosts in the following order (in decreasing preference): recycling hosts whose lifetime classes are greater than the VM lifetime class (the closer the more preferred), open hosts whose lifetime class are equal to the VM lifetime class, any non-empty host, and lastly empty hosts. In contrast to NILAS, which acts as a tie-breaker for the existing algorithm, LAVA makes decisions that the baseline could not have made, and is therefore invasive.
@@ 4.4 LARS: Defragmentation & Maintenance
The above algorithms focus on reducing fragmentation by improving VM placement. Borg also performs software and hardware maintenance activities, applying security, microcode, host OS, and VM software updates. It also actively defragments hosts for some VM families. These activities live migrate VMs Ruprecht et al. (2018). Live migration uses pre- and post-copy techniques that minimize VM execution time disruptions to seconds. However, total migration times are proportional to VM memory and SSD usage, and consume capacity on both hosts. In simulations, we model that both hosts are busy for a conservative 20 minutes.
Borg reserves a small fraction of hosts for defragmentation, maintenance, and to tolerate hardware failures, and thus these activities do not impact available capacity. We show how using lifetimes can reduce the number of VMs that are disrupted during defragmentation and maintenance. We call the algorithm LARS. The rest of our exposition focuses on defragmentation for conciseness (details in Appendix H).
@@
Note that LARS uses the same algorithm to choose a target host for migration as for initial VM placement. As such, NILAS and LAVA use repredictions from our distribution-based model to place the VM on a better-suited host compared to the prior algorithm that was not lifetime-aware.
@@ 5 Implementation
Our algorithms are integrated into Borg’s scheduling code. The scheduler runs within Borg Prime, which handles incoming VM requests and assigns them to hosts. The Borg Prime binary is replicated across a number of instances, with one of them serving as the elected leader. All replicas run the same binary and a lower layer implements consensus.
As mentioned in Section 3, we compile our model into the Borg Prime binary and roll it out with Borg. This coupling has the advantage that canary deployments (the first step of our rollout process) test the entire end-to-end approach. We deploy our software regularly and could update our model as often as once a week. However, we found that the models’ accuracy remains high even a month later as shown in Figure 10. Note that while we found it highly advantageous from a latency and complexity perspective to compile the model into the binary, our approach is not restricted to this deployment scenario.
We carefully measured the CPU and DRAM overheads of our approach in production. Binary size increases by less than 10 MB, and both the memory and CPU consumption of our approach are (generally) indistinguishable from noise. This lack of observable overhead may seem counter-intuitive, but since a VM record has so much other state and Borg performs many other computation during scoring, the extra bytes of memory and the 9 us of model latency are not visible. We process on the order of 10-100 scheduling requests per second in each cluster, and even hypothetically repredicting 100 VMs in the process would add less than 1 ms. In practice, we generally repredict many fewer VMs, since higher-ranking scoring functions filter out many hosts and we re-score in parallel VMs only on considered hosts.
In some very large zones, re-scoring still became a bottleneck. We therefore added a host lifetime score cache (Appendix G.3). For any host that 1) is assigned a new VM, or 2) an existing VM exits, or 3) the host expected lifetime expires, we re-score the host. For all other hosts, we re-score them according to a configurable time interval (e.g., 60s).
@@ 5.1 Simulations
NILAS/LAVA: We developed an event-driven simulator to explore our approach. We extract production traces of VM start, exit, and restart events within a particular pool and then replay this trace against a simulated instance of the scheduler. The simulated scheduler uses our production code base and is thus highly accurate. We validated our simulator against production (Appendix F has more details).
LARS: To calculate the benefit of LARS (lifetime-based defragmentation), we built a modified simulator, using the same production traces. From our traces, we collect the list of migrations that were performed during a time interval. In production, we limit the number of live migrations that can be in progress simultaneously to batches of 3. We simulate this behavior by assuming that all migrations are performed in a certain order (in our baseline, defined by the trace), but have to wait until a slot is available. This approach has the effect that some VMs exit while others are migrating. LARS modifies this order based on lifetime predictions.
@@ 5.2 Production Measurements
Measuring the impact of scheduling policies in production is challenging. Since workloads and total capacity in a pool are constantly changing, it is difficult to attribute changes in a metric to a particular modification. We therefore use an A/B testing methodology. We split the hosts into two halves and apply our scheduling algorithm to one of them. Our production metrics report empty hosts, stranding, and model prediction for the entire pool. However, stranding is only accurate for entire pools and cannot differentiate A/B hosts. We report live production experiments in four pools, where we apply the algorithm to the entire pool and evaluate scheduling quality. These experiments also ensure that our A/B approach does not distort behavior. Note that these A/B deployments are at production scale. The cluster/zone sizes for this product range from 100s to 10,000+ hosts with O(100-10,000) VMs active at a time.
NILAS has been deployed fleet-wide for one VM family for about a year, at the time of publication. Time series data from the fleet-wide production rollout confirmed our simulation numbers and A/B experiments, and show a fleet-wide increase of empty hosts by 4 pp for this VM family.
@@ 5.3 LA-Binary Comparison
Our main comparison is to LA-Binary, a faithful implementation of LA, but embedded in the Borg binary versus a separate model service Barbalho et al. (2023). Since our data center environment is different, perfectly replicating their setup is impossible as, e.g., the ML model features are different. We are faithful to LA’s description and favorable to LA-Binary when in doubt. For an apples-to-apples comparison, we use the exact same model when comparing LA-Binary to NILAS and LAVA, but without repredictions for LA-Binary. The predictor classifies VMs as short or long lived, using all our model features, but with the same two-hour cutoff threshold as in their paper.
Figure 6: Empty host improvements comparing NILAS, LAVA with LA-Binary, to our production baseline for 24 C2 pools in simulation. On average, LAVA increases empty hosts by 6.5 pp, compared to 6.1 pp for NILAS, and 5.0 pp for LA-Binary.